OpenStudy (anonymous):

the line passes through the orgin and has a slope of 4. what is the equation of the line that is parallel to the first line and passes through the point (1,-2)

OpenStudy (dls):

(0,0) and m=4 Use \[\Huge y-y_1=m(x-x_1)\] y1=0 x1=0

OpenStudy (dls):

Similarly you have points and m in 2nd case as well

OpenStudy (anonymous):

Let the equation of the original line be y=mx+c where m is the slope of the line and(x,y) is the required point.... now since the line passes through origin....so origin satisfies the equation of the line... 0=m(0)+c c=0 thus we get y=mx in this case m=4 thus eqution of original line is y=4x ......... now we have the find the equation of such a line which is parallel to the above line .....(parallel lines have same slope) so the slope of this line is also equal to 4 secondly it passes through (1,-2)...therefore -2=4(1)+c -2-4=c c=-6 thus the final equation of the new line is y=4x-6...(in slope form)

OpenStudy (dls):

You are not supposed to give out the answer but guide the asker to it.

OpenStudy (anonymous):

sorry.....

OpenStudy (anonymous):

thanks nitz anyway so much

OpenStudy (anonymous):

u are welcome..